# An Adverse Weather-Immune Scheme with Unfolded Regularization and Foundation Model Knowledge Distillation for Street Scene Understanding

### v1

@@ Abstract
Various adverse weather conditions pose a significant challenge to autonomous driving (AD) perception. A common strategy is to minimize the disparity between images captured in clear and adverse weather conditions. However, this technique typically relies on utilizing clear image as a reference, which is challenging to obtain in practice. Furthermore, this method typically targets a single adverse condition and perform poorly when confronting the mixup of multiple adverse weather conditions. To address these issues, we introduce a reference-free and Adverse weather-Immune scheme (called AdvImmu) achieved by leveraging the invariance of weather conditions over short periods (seconds). Specifically, AdvImmu includes three components: Locally Sequential Mechanism (LSM), Globally Shuffled Mechanism (GSM), and Unfolded Regularizers (URs). LSM leverages temporal correlations between adjacent frames to enhance model performance. GSM is proposed to shuffle LSM segments to prevent the overfitting to temporal patterns of only using LSM. URs are the deep unfolding implementation of two proposed regularizers to penalize the model complexity to enhance across-weather generalization. In addition, to overcome the over-reliance on consecutive frame-wise annotations in the training of AdvImmu (typically unavailable in AD scenarios), we incorporate the Segment Anything Model (SAM) to annotate frames, and additionally propose a cluster algorithm (denoted as SBICAC) to surmount SAM’s category-agnostic issue to generate pseudo-labels. Extensive experiments demonstrate that the proposed AdvImmu outperforms existing state-of-the-art methods by 88.56% in mean Intersection over Union (mIoU).
@@ I Introduction
Autonomous driving (AD) perception is critical [1, 2, 3] and is still a significant challenge in various adverse weather conditions so far [4]. Existing domain adaption approach to address this challenge involves reducing the disparity between images captured in clear weather as a reference and those taken in adverse conditions [5, 6]. However, obtaining such a reference image can be challenging in reality. Furthermore, these methods are generally developed to handle specific adverse condition, such as foggy weather. This specialization often leads to a significant decline in performance when vehicles encounter unusual weather conditions where the model has not been specifically tuned.
@@
To effectively tackle these challenges, we propose a novel scheme (named AdvImmu). AdvImmu benefits from being independent of reference images and is robust against various adverse weather conditions, which is materialized by considering the stability of weather conditions within past seconds. Concretely, AdvImmu consists of Locally Sequential Mechanism (LSM), Globally Shuffled Mechanism (GSM), and Unfolded Regularizers (URs) modules. Specifically, LSM exploits the temporal correlations between consecutive frames via three distinct units: Instant Unit (InsU), which is designed to detect instantaneous information in the input; Integral Unit (IntU), which focuses on capturing stable and shared background information from the scenes; Derivative Unit (DU), which excels in sensing dynamic changes in the environment. As discussed, since LSM integrates temporal patterns from historical consecutive frames, it is susceptible to overfitting specific input sequence order. To mitigate LSM’s temporality-related overfitting, GSM is introduced to rearrange the segments processed by the LSM. This strategic shuffling ensures that the model does not merely memorize a particular data order but instead learns a generalizable pattern across varied data sequences. On top of LSM and GSM, URs are proposed to improve the model’s generalization across diverse adverse weather conditions by penalizing the model complexity. Concretely, in addition to the conventional pixel-level cross entropy loss used in semantic understanding, we propose using the deep unfolding technique [7] to unfold an image-level regularizer and an inter-class contrastive regularizer into layers. Deep unfolding can avoid heuristics or exhaustive searches on the weights of both regularizers, which prevents instability and suboptimal performance. The proposed AdvImmu is summarized in Fig. 1.
Figure 1: Overview of the proposed AdvImmu. GSM shuffles and groups the LSM segments to avoid the overfitting to temporal patterns, whereas GSM-Free does not. In the context of LSM, $T$ denotes the current frame number in the input sequence, while $t$ (i.e., LSM depth) refers to the count of consecutive frames prior to frame $T$.
The training of the proposed AdvImmu requires consecutive frame-wise annotation for the images in training datasets, which is impractical in the realm of AD scenarios. To tackle this challenge, we propose to use the Segment Anything Model (SAM) [8] to annotate the images. However, since SAM-generated masks lack of category information (e.g., different trees are assigned with different semantic IDs), we additionally propose a Similarity-based Iterative Cluster Assignments and Centroids Refinement (SBICAC) algorithm to overcome SAM’s category-agnostic issue and further to assign unique ID to each semantic class as frame-wise labels. Based on SAM and SBICAC, we can train the proposed AdvImmu in an unsupervised manner.
The main contributions are summarized as follows:
@@
  • The training of AdvImmu needs to overcome SAM’s category-agnostic issue in the output. To this end, SBICAC algorithm is proposed to generate category IDs serving as training labels based on SAM’s output.
@@ II Related Work
II-A AD Perception under Adverse Weather Conditions
AD perception under various adverse conditions has recently attracted increasing attention due to the strict safety demand in various outdoor navigation [9, 10] and perception tasks [11, 12, 13, 14]. Domain adaption [15] is the major scheme in enhancing AD perception capabilities under various adverse conditions. Some studies in domain adaption [6, 16] aim to learn a generalizable model from single or multiple related but distinct source domains where target data is inaccessible during model learning. Some adversarial training-based domain adaption works attempt to narrow the domain gap via style transfer or learning indistinguishable representations [17]. Furthermore, equipped with abundant data from various domains, curriculum learning-based domain adaption realize a progressive adaptation towards a distant target domain [18]. In addition, extending domain adaptation to multiple target domains usually employs domain transfer [19] to enhance generalization across domains. In this work, unlike domain adaption, the proposed AdvImmu can sense instantaneous, stable backgound, and dynamic information by considering temporal correlations, which is robust to arbitrary adverse conditions.
@@ II-B Large Models (LMs) in AD
Large Models (LMs) have gained significant interest in the context of AD due to their proficiency in analyzing images and contexts [20]. These advancements have greatly improved zero-shot and few-shot image classification [21], segmentation [22, 23], object detection [24], etc. For example, CLIP [25] have shown that training to match images with captions can effectively create image representations from scratch; LLaMa [26] combines a vision encoder with an Large Language Model to enhance the understanding of both visual and linguistic concepts. Furthermore, researchers have explored the vectorized visual embeddings to equip LMs with environmental perception capabilities, particularly in AD scenarios. For example, DriveGPT4 [27] interprets video inputs to generate driving-related textual responses; Talk2BEV [28] leverages pre-trained image-language models to combine Bird’s Eye View maps with linguistic context, enabling visuo-linguistic reasoning in autonomous vehicles. In addition, generative models in AD are also gaining popularity. For instance, GAIA-1 [29] generates realistic driving scenarios by integrating video, text, and action inputs and showcases generative models in adapting to the changing dynamics of the real world. In this work, we propose to use SAM [8] to generate frame-wise annotation. However, SAM’s output lacks of category information. To mitigate this limitation, we additionally propose SBICAC clustering algorithm to supplement SAM to generate category IDs.
@@ III Methodology
To depict the proposed AdvImmu clearly, we will introduce LSM and GSM together and denote the combination of them as LSaGS hereafter. Subsequently, we will detail URs and how the proposed SBICAC algorithm along with SAM to generate pseudo-labels.
@@ III-A LSaGS
LSM extracts immediate, background, and dynamic features through InsU, IntU, and DU, respectively. Concretely, InsU is designed to process the current image as vehicle drives. It extracts instantaneous information from this image, which is crucial for the vehicle to accurately understand immediate surroundings. IntU focuses on processing stable background information from the environment, achieved by analyzing past consecutive images prior to the current frame. This enables IntU to understand what remains constant over time, which is important for consistent understanding of the scene. DU analyzes dynamic changes in the environment by comparing the current frame against previous ones, which helps to highlight the dynamic objects, such as pedestrians and vehicles. Notably, LSM depth is used to indicate the count of past consecutive frames that are considered.
@@
Early Fusion (EF) and Late Fusion (LF) are two policies for fusing features from InsU, IntU, and DU. In EF strategy, features from InsU, IntU, and DU are first merged. Then the merged features are processed by a shared Convolution Neural Network (CNN) model. This early integration allows the model to utilize the correlations between these different types of features, enhancing the model for understanding street scenes. On the other hand, in LF strategy, InsU, IntU, and DU process the their own features by three separate CNN models. This allows each type of features to be finely analyzed on its own. After processing, these different types of features are combined. EF is summarized in the right part of Fig. 1 and LF is outlined in Fig. 2.
@@
GSM is introduced to shuffle and group different LSM segments into mini-batches. In this way, each mini-batch reflects the diversity of the dataset instead of following a strict sequence. This enables the model learn more general and robust patterns, enhances its performance and generalization. Based on the shuffled mini-batches by GSM, the involved CNN(s) and the downstream perception head are then optimized by back propagation.
In summary, LSaGS is finally outlined in Algo. 1.
@@
7                   $MF_{t}\leftarrow\text{Merge}(InsU_{t},IntU_{t},DU_{t})$
8                   $PF_{t}\leftarrow\text{CNN}_{\text{shd}}(MF_{t})$

### v3

# Generalizable Autonomous Driving System across Diverse Adverse Weather Conditions

@@ Abstract
Various adverse weather conditions pose a significant challenge to autonomous driving (AD) street scene semantic understanding (segmentation). A common strategy is to minimize the disparity between images captured in clear and adverse weather conditions. However, this technique typically relies on utilizing clear image as a reference, which is challenging to obtain in practice. Furthermore, this method typically targets a single adverse condition, and thus perform poorly when confronting a mixture of multiple adverse weather conditions. To address these issues, we introduce a reference-free and Adverse weather-Immune scheme (called AdvImmu) that leverages the invariance of weather conditions over short periods (seconds). Specifically, AdvImmu includes three components: Locally Sequential Mechanism (LSM), Globally Shuffled Mechanism (GSM), and Unfolded Regularizers (URs). LSM leverages temporal correlations between adjacent frames to enhance model performance. GSM is proposed to shuffle LSM segments to prevent overfitting of temporal patterns. URs are the deep unfolding implementation of two proposed regularizers to penalize the model complexity to enhance across-weather generalization. In addition, to overcome the over-reliance on consecutive frame-wise annotations in the training of AdvImmu (typically unavailable in AD scenarios), we incorporate a foundation model named Segment Anything Model (SAM) to assist to annotate frames, and additionally propose a cluster algorithm (denoted as SBICAC) to surmount SAM’s category-agnostic issue to generate pseudo-labels. Extensive experiments demonstrate that the proposed AdvImmu outperforms existing state-of-the-art methods by 88.56% in mean Intersection over Union (mIoU).
Index Terms:
Semantic Segmentation, Street Scene Understanding, Mixup of Adverse Weather Conditions, Unfolded Regularization, Segment Anything Model (SAM), Knowledge Distillation.
@@ I Introduction
Autonomous driving (AD) perception is critical [1, 2, 3] for AD vehicles. In general, AD perception consists of a set of tasks, such as semantic segmentation [4, 5], object detection [6], object tracking [7], etc. This work stays focused on semantic segmentation which aims to classify each pixel into predefined categories. Currently, semantic segmentation models are typically built using Deep Learning (DL) techniques [8, 9, 10, 11, 12], particularly Convolutional Neural Networks (CNNs) [13, 14] and Transformers [15]. These models are generally trained on large datasets with labels that are manually annotated at the pixel level. Once trained, these models can generally predict the category of each pixel of given images properly.
Figure 1: Major challenges in the realm of domain adaption [16]. (a) Requirement of reference images. Reference images, highlighted in the red box located at the bottom right corner, depict clear-weather-condition scenes which are almost same with those under adverse weather conditions (e.g., nighttime or fog). Generally, it is difficult to collect such reference images in practice. (b) Poor performance under the hybrid of multiple adverse conditions. In general, domain adaption models always work well in adapted target domains and face significant performance decline in unadapted domains, which results in poor performance when dealing with the case of hybrid of multiple adverse weather conditions.
@@
Although such DL-based models achieve advanced performance for semantic segmentation task, they generally face significant performance drop in various adverse weather conditions [17]. For example, a semantic segmentation model trained on a dataset predominantly composed of clear, sunny weather conditions, will decay significantly when the vehicle encounters a foggy environment. This is because the fog causes a reduction in visibility and alters the appearance of objects, leading to incorrect prediction of each pixel’s category. Existing domain adaption approach to address this challenge involves reducing the disparity between images captured in clear weather as a reference and those taken in adverse conditions [16]. However, obtaining such a reference image can be challenging in reality. Furthermore, these methods are generally developed to handle specific single adverse condition, such as foggy weather. This specialization often leads to a significant decline in performance when vehicles encounter other adverse weather conditions where the model has not been specifically tuned. The mentioned challenges are summarized in Fig. 1.
Figure 2: Overview of the proposed AdvImmu. In the context of LSM, Instant Unit (InsU), Integral Unit (IntU), and Derivative Unit (DU) capture instantaneous information, stable and shared background information, and dynamic changes from the input, respectively. These three units work together to enhance model performance by leveraging temporal correlation of intra-weather condition. Notably, $T$ denotes the current frame number in the input sequence, while $t$ (i.e., LSM depth) refers to the count of consecutive frames prior to frame $T$. GSM shuffles and groups the LSM segments to avoid the overfitting to specific temporal patterns, whereas GSM-Free does not. GSM is proposed to enable the model to learn common patterns across multiple adverse weather conditions. URs include two unfolded regularizers and are proposed to penalize the model complexity to avoid overfitting to specific patterns as well, therefore, enhancing the model performance across multiple adverse weather conditions.
To effectively tackle these challenges, we propose a novel scheme (named AdvImmu). AdvImmu benefits from being independent of reference images and is robust against a mixture of various adverse weather conditions. The proposed method is materialized by considering the local temporal correlation, global randomness, and unfolded regularization. Specifically, AdvImmu consists of Locally Sequential Mechanism (LSM), Globally Shuffled Mechanism (GSM), and Unfolded Regularizers (URs) modules.
@@
In particular, LSM exploits the temporal correlations between consecutive frames via three distinct units: Instant Unit (InsU), which is designed to detect instantaneous information in the input; Integral Unit (IntU), which focuses on capturing stable and shared background information from the scenes; Derivative Unit (DU), which excels in sensing dynamic changes in the environment. However, since LSM integrates temporal patterns from historical consecutive frames, it is susceptible to overfitting specific input sequence order. To mitigate LSM’s temporality-related overfitting, GSM is introduced to rearrange the segments processed by the LSM. This strategic shuffling ensures that the model does not merely memorize a particular data order but instead learns a generalizable pattern across various data sequences. On top of LSM and GSM, URs are proposed to improve the model’s generalization across diverse adverse weather conditions by penalizing the model complexity. Concretely, in addition to the conventional pixel-level cross entropy loss used in semantic understanding, we propose to use the deep unfolding technique [18] to unfold an image-level regularizer and an inter-class contrastive regularizer into layers. Deep unfolding can avoid heuristic or exhaustive searches on the weights of both regularizers, which prevents instability and suboptimal performance. The proposed AdvImmu is summarized in Fig. 2.
In general, the training of the proposed AdvImmu requires consecutive frame-wise annotations for all images in the training sequences. As it is impractical to annotate manually in the realm of AD scenarios, we propose to use a foundation model named Segment Anything Model (SAM) [19] to assist annotating images in the training sequences, thanks to SAM’s powerful capabilities to segment objects [20]. However, SAM-generated masks lack category information (demonstrated in LABEL:Fig.SAM_reason_to_cluster). To overcome this category-agnostic issue, [21] propose to use prompts to assist SAM to generate precise class masks. This method is difficult to automatically process batches of images owing to the requirement of articulated prompts. [22] utilizes Class Activation Maps (CAM) pseudo-labels as cues to select and combine SAM masks to generate class-aware pseudo-labels, which, however, needs extra CAM-related classification model and post-processing techniques. Therefore, this method is also impractical for generating consecutive frame-wise annotations.
In this work, we propose a novel clustering algorithm to supplement SAM to generate category-aware pseudo-labels. Specifically, we propose a clustering algorithm named Similarity-based Iterative Cluster Assignments and Centroids Refinement (SBICAC) to overcome SAM’s category-agnostic issue and further assign unique ID to each semantic class as frame-wise labels. SBICAC performs clustering on SAM-extracted latent feature maps. The process begins by initializing cluster centroids randomly. It then iterates through a predefined maximum number of iterations. In each iteration, the proposed SBICAC: (I) Computes Similarity: It calculates the similarity between each pixel’s latent feature and all centroids using a similarity metric (e.g., inner product or cosine similarity). This operation results in a similarity matrix. (II) Assigns Clusters: Each pixel is assigned to the cluster with the highest similarity, thus updating the labels for each pixel based on maximum similarity. (III) Updates Centroids: For each cluster, the new centroid is recalculated as the mean features of all pixels assigned to that cluster. (IV) Checks for Convergence: The loop checks if the centroids have stabilized (i.e., the old centroids and new centroids are sufficiently close to each other). If they have converged, the loop breaks, ending the iterations early. The output of the algorithm is the cluster assignments (labels) for all pixels. In summary, SBICAC iteratively refines pixels’ cluster assignments and centroids based on similarity until the convergence is achieved. Based on the high-quality pseudo-labels generated by SBICAC and SAM, we can train the proposed AdvImmu in an unsupervised manner.
The main contributions of this paper are highlighted as follows:
@@
  • The training of AdvImmu needs to overcome SAM’s category-agnostic issue in the output. To this end, an unsupervised clustering algorithm (i.e., SBICAC) is proposed to generate category IDs serving as training labels based on SAM’s output.
@@
The remainder of this paper is organized as follows. Section II provides an overview of the related works. Section III elaborates on the proposed methodologies. Section IV presents a comprehensive set of experiments along with empirical analyses. Section V concludes this paper.
@@ II Related Work
II-A Domain Adaption for Autonomous Driving Perception under Adverse Weather Conditions
AD street scene semantic understanding under various adverse conditions has recently attracted increasing attention due to the strict safety demand in various outdoor perception [23, 24] and navigation [25] tasks. Domain adaption [26] is the major scheme in enhancing AD capabilities of street scene semantic understanding under various adverse conditions. Studies in domain adaption [27] aim to learn a generalizable model from single or multiple related but distinct source domains where target data is inaccessible during model learning. Most domain adaption approaches can be grouped into adversarial-based domain adaption, curriculum learning-based domain adaption, and self-training-based domain adaption.
Specifically, adversarial training-based methods for domain adaptation aim to minimize the domain discrepancy through style transfer or by learning representations that are indistinguishable across domains [28]. Typically, these methods work within a GAN framework [29] to align the distributions of the source and target domains at various levels: input [30], feature [31], or output [32]. On the other hand, curriculum learning-based domain adaptation achieves gradual adaptation to a more distant target domain [33] by utilizing extensive data from multiple domains. Finally, self-training-based domain adaptation involves adapting the model using high-confidence pseudo-labels generated from the target domain data. To regulate training and prevent drift in pseudo-labels, strategies such as confidence thresholding [34], pseudo-label prototypes [35], and consistency regularization through data augmentation [36] and varying contexts [37] are employed. Moreover, when extending domain adaptation to encompass multiple target domains, domain transfer techniques are applied [38] to improve generalization across these domains.
Domain datasets frequently exhibit a long-tail distribution, leading to a model bias towards more prevalent classes [39]. In the field of domain adaptation, strategies such as loss re-weighting [34] and class-balanced sampling tailored for image classification [40] have been implemented to address this issue. HRDA [15] introduces Rare Class Sampling (RCS) to extend these techniques from classification tasks to semantic segmentation, which tackles the challenge of both rare and common classes appearing within each semantic segmentation category. HRDA [15] has also expanded the application of scale attention [41] to domain adaptation, highlighting its critical role in enhancing domain robustness by improving the generalization capabilities across varying object sizes.
Some semantic segmentation models often incorporate multi-scale features, a practice popularized by the widespread use of DeepLabV2 in domain adaptation [42]. Subsequent studies have explored techniques such as multi-scale average pooling during inference [36] and enforcing scale consistency [43, 44] for low-resolution inputs to facilitate multi-scale feature fusion. In addition, previous research [45] has demonstrated that using knowledge distillation from prior tasks can serve as a regularization mechanism for new tasks, thereby enhancing the performance of domain adaptation [46].
In this work, unlike domain adaption, the proposed AdvImmu can sense instantaneous, stable background, and dynamic information by considering temporal correlations, which is robust to arbitrary adverse conditions.
@@ II-B Large Models (LMs) in AD
Large Models (LMs) have gained significant interest in the context of AD due to their proficiency in analyzing images and contexts [47]. These advancements have greatly improved zero-shot and few-shot image classification [48], segmentation [49], object detection [50], etc. For example, CLIP [51] have shown that training to match images with captions can effectively create image representations from scratch; LLaMa [52] combines a vision encoder with an Large Language Model to enhance the understanding of both visual and linguistic concepts.
Furthermore, researchers have explored the vectorized visual embeddings to equip LMs with environmental perception capabilities, particularly in AD scenarios. For example, DriveGPT4 [53] interprets video inputs to generate driving-related textual responses; Talk2BEV [54] leverages pre-trained image-language models to combine Bird’s Eye View maps with linguistic context, enabling visuo-linguistic reasoning in autonomous vehicles.
Generative models in AD are also gaining popularity. For instance, GAIA-1 [55] generates realistic driving scenarios by integrating video, text, and action inputs and showcases generative models in adapting to the changing dynamics of the real world. In this work, we propose to use SAM [19] to assist to generate frame-wise annotation. However, SAM’s output lacks of category information. To mitigate this limitation, we additionally propose SBICAC clustering algorithm to supplement SAM to generate category-aware annotations.
II-C Knowledge Distillation
In the domain of DL, knowledge distillation is defined as the technique where knowledge is transferred from a large, intricate model (i.e., teacher model) to a smaller, more streamlined model (i.e., student model) [56]. This method is crucial for addressing the computational and resource limitations associated with the deployment of large-scale models in real-world settings.
Prior to the advent of LMs, knowledge distillation primarily focused on the transfer of knowledge from a sophisticated and often large teacher model to smaller and more efficient student models [57]. The drive behind this process was the necessity to implement machine learning models in environments constrained by computational and memory resources, such as mobile or edge computing devices. The emphasis was mainly on selecting neural architectures and designing training objectives specifically for individual tasks. Earlier approaches included training a compact student model to replicate the output the teacher model. Generally, employing methods like soft target training to enable the student model to learn from the softened softmax outputs of the teacher model.
With the emergence of LMs, the focus of knowledge distillation has significantly evolved. Unlike previous approaches that primarily targeted architectural compression, the modern approach in LM era emphasizes extracting and transferring knowledge [58]. This shift is driven by the rich and intricate knowledge embedded within LMs such as SAM [19]. Moreover, the vast and often unmanageable number of parameters in LMs complicates the use of traditional size-reduction techniques like pruning [59] or quantization [60]. Currently, the objective of knowledge distillation in the context of LMs is not merely to replicate the teacher model’s output or to reduce the model size, but rather to harness and transfer the distinct knowledge contained within these powerful LMs.
The modern approach to knowledge distillation within LMs centers around the use of crafted prompts. These prompts are designed to specifically elicit targeted knowledge [61] or capabilities [58] from LMs. They effectively access the LMs’ comprehension and skills across a spectrum of areas, from natural language understanding [62] to more sophisticated cognitive functions such as reasoning [63] and problem-solving [64]. The strategic use of prompts introduces a effective and flexible method of knowledge extraction, enabling precise targeting of skills or areas of interest. This technique proves especially beneficial in leveraging the emergent abilities of LMs.
Additionally, the new paradigm in knowledge distillation extends mere output replication to transfer of more nuanced attributes like reasoning patterns [65], preference alignment [66], and value alignment [67]. This shift represents a move towards a broader and more in-depth transfer of cognitive skills. Current practices not only replicate outputs but also emulate the cognitive processes [65] and decision-making approaches [68] of the teacher model. Techniques such as chain-of-thought (CoT) prompting are employed, where the student model is trained to internalize and replicate the teacher’s reasoning pathway, thus enhancing its capabilities in problem-solving and decision-making. In this paper, to generate consecutive frame-wise annotations, we propose to distill knowledge from SAM [19] and then use the distilled knowledge along with the proposed unsupervised SBICAC clustering algorithm to generate pseudo-labels for consecutive frames.
@@ III Methodology
To depict the proposed AdvImmu clearly, we denote the combination of LSM and GSM as LSaGS hereafter. We will introduce LSaGS in Section III-A. Subsequently, we will detail URs in Section III-B, including the definition of both regularizers and the proposed unfolding method. Finally, we will present how the proposed SBICAC algorithm along with SAM to generate frame-wise pseudo-labels in Section III-C.
@@ III-A LSaGS
LSM is proposed to process local temporal dependency to enhance the model performance. Specifically, LSM extracts immediate, stable background, and dynamic features through InsU, IntU, and DU, respectively. Concretely, InsU is designed to process the current image as the vehicle drives. It extracts instantaneous information from this image, which is crucial for the vehicle to accurately understand immediate surroundings. IntU focuses on processing stable background information from the environment, achieved by analyzing past consecutive images prior to the current frame. This enables IntU to understand what remains constant over time, which is important for consistent understanding of the scene. DU analyzes dynamic changes in the environment by comparing the current frame against previous ones, which helps to highlight the dynamic objects, such as pedestrians and vehicles. In this work, we propose to use ResNet50 [69] to serve as the network architecture of InsU, IntU and DU. Notably, LSM contains a hyper-parameter named LSM depth, which is used to indicate the count of the considered consecutive frames prior to current frame. LSM depth definitely influences IntU and DU.
@@
Early Fusion (EF) and Late Fusion (LF) are two policies for fusing features from InsU, IntU, and DU. In EF strategy, features from InsU, IntU, and DU are first merged. Then the merged features are processed by a shared CNN model. This early integration allows the model to utilize the correlations between these different types of features, enhancing the model for understanding street scenes. On the other hand, in LF strategy, InsU, IntU, and DU process their own features by three separate CNN models. This allows each type of features to be finely analyzed on its own. After processing, these different types of features are combined. EF is summarized in the right part of Fig. 2 and LF is outlined in Fig. 4.
@@
GSM is introduced to shuffle and group different LSM segments into mini-batches. In this way, each mini-batch can reflect the diversity of the dataset instead of following a strict sequence. This enables the model learn more general and robust patterns, enhances its performance and generalization. Based on the shuffled mini-batches by GSM, the involved CNN models are then optimized by back propagation.
After the introduction of the philosophy of LSM and GSM, we then formulate LSaGS in Algorithm 1 as a whole picture of LSaGS. Specifically, LSaGS iterates multiple rounds. In each round, it follows below steps: (I) Extract features: LSaGS takes video sequence $F$ as input to process. For frame $F_{t}\in F$, InsU (with parameters $\phi_{\text{InsU}}$) processes immediate features $InsU_{t}$ from $F_{t}$. IntU (with parameters $\phi_{\text{IntU}}$) analyzes the background by looking at the past $d$ frames (i.e., $F_{t-i}$ for $i=1$ to $d$) to gather historical context features $IntU_{t}$. DU (with parameters $\phi_{\text{DU}}$) examines changes in the current frame $F_{t}$ relative to the past $d$ frames, capturing dynamic change features $DU_{t}$. (II) Merge features: After extracting features by InsU, IntU, and DU, LSaGS performs feature fusion based on the specified fusion method. For EF, $InsU_{t}$, $IntU_{t}$, and $DU_{t}$ are merged into $MF_{t}$, which is then processed by a shared CNN (with parameters $\text{CNN}_{\text{shd}}$) to output the processed features $PF_{t}$. For LF, $InsU_{t}$, $IntU_{t}$, and $DU_{t}$ are processed separately by their respective CNNs (with parameters $\text{CNN}_{\text{InsU}}$, $\text{CNN}_{\text{IntU}}$, $\text{CNN}_{\text{DU}}$, respectively) to extract features $PInsU_{t}$, $PIntU_{t}$ and $PDU_{t}$. Such features are then merged to form the final processed features $PF_{t}$. (III) Group and Shuffle Mini-Batches: After processing all frames, LSaGS shuffles and groups $PF_{t}$ into mini-batches. This step helps in preparing the data for training. (IV) Optimization via back propagation: LSaGS is optimized by back propagation in a batch-by-batch way.
@@
7                   $MF_{t}\leftarrow\text{Merge}(InsU_{t},IntU_{t},DU_{t})$;
8                   $PF_{t}\leftarrow\text{CNN}_{\text{shd}}(MF_{t})$;